From nch@roe.ac.uk Fri Feb 20 13:31:40 2004 Date: Fri, 20 Feb 2004 12:04:01 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Harvey MacGillivray , Ian Bond , Mike Read , Nigel Hambly , Bob Mann Cc: CCs for WSA weekly meeting minutes distribution -- Clive Davenhall , Andrew Lawrence , Andy Adamson , Peter Shillan , John Taylor , Jim Emerson , Martin Hill , Mike Irwin , Peredur Williams Subject: WFAU WSA weekly meeting minutes: 20th February 2004 Minutes of WFCAM Science Archive meeting: 20th February 2004 ------------------------------------------------------------ ------------------------------------------------------------ Present: MAR, JMS, GPS, MCH, NCH, ETWS, PMW, IAB, JDT Apologies: JPE, AL, RGM, HMG DONM: 10am, Friday 27th February 2004, plate library. Actions discharged: ------------------- ACTION: NCH to chase up Eclipse concerning the status of Tyan motherboard U320 controller firmware updates. Done - NCH will now liaise with JNTD as to testing the latest updates which Eclipse have sent. ACTION: IAB to contact MJI/JRL to ask about the possibility of some more example calibration frames (eg. darks, flats, skys, defringes, object masks). Done - see Software. Actions partly discharged but continuing: ----------------------------------------- ACTION: ALL to review risk register and think of any new internal or external risks that should be documented in the risk register. Defered until top level plan is explicitly defined. PMW tabled the VISTA risk register as an example. ACTION: JPE to design and set up centralised VDFS web pages. Progressing. (there is now a link to it from our WSA Twiki). - continuing; thanks for progressing these. Actions carried forward from 23/01/04 meeting: ----------------------------------------------- None. Specific points and new actions: -------------------------------- Project management: NCH noted that JPE had sent around a drafty VDFS User Requirements Document; NCH has sent comments to JPE and WJS. JMS noted that the origin of the doc is in the documentation set for ESO deliverables and hence it's emphasis on pipeline processing and calibration; NCH noted that any requirements over and above those for the WSA must be documented for the purposes of the reviews (PDR and FDR) currently scheduled in the overall VDFS plan. Comments and issues arising from CASU fortnightly minutes: Nothing new this week. Networking: Nothing new this week. Hardware: NCH chased up Eclipse concerning Tyan motherboard firmware updates, and we have received software updates yesterday. NCH & JNTD will now try these on server amenhotep to check if the U320 problem is solved. NCH noted that apart from the occaisional glitch, both servers appear to be performing well as regards their disks. NCH noted that the aircon in C1 now appears to be performing well. Software: IAB has asked MJI & JRL for some example calibration frames to test the ingest procedure; the CASU folks are now looking into this. IAB reported: "Have implemented the frame ID method using the prescription as discussed. Standalone code for CVS generation is ready for integration into the curation scripts. It takes a few minutes to generate CSV files from and input of ~1000 FITS files." NCH thanked ETWS for "spring cleaning" the CVS (ie removing some deprecated and/or empty directories). NCH & ETWS have been starting to think about the procedure for CU19 and CU20 (survey DB freeze, backing up, release, place on line) and NCH noted that there are some great tailor-made features in SQL Server to facilitate this (namely "data replication"). NCH requested that MAR & ETWS stayed back after the meeting to look over the detail of this. IAB & NCH have continued to work on the CU3 and CU4; another integration session is pencilled in for Monday ACTION: IAB & NCH to meet Monday pm for another integration session on the CU3 and CU4 software. NCH noted that the WSA external catalogues (2MASS, ROSAT, USNOB, IRAS, FIRST, and SDSS DR1) are now loaded, and intermediate ingest files and database files backup on LTO-2. This frees up wiglaf (which was being used as a temporary staging area for some of the intermediate files). Some discussion ensued concerning whether to keep the old version of SDSS EDR (ie. the flat files associated with the SX Objy catalogue server). MAR suggested switching off the interface to see if anyone cries foul. ACTION: ETWS to switch off web access to the wiglaf SDSS SX EDR flat file system. SSA: NCH noted that the SSA Source table has completely reloaded with the new galactic co-ord/extinction attributes; some performance problems have been encountered while reattaching foreign key constraints - NCH is looking into this. Otherwise, once the firmware updates have been tested and assuming the disk IO problems are solved, the terabyte SSA server will be ready to be place online. Miscellaneous: NCH noted that the VOProcPlus meeting in Leicester on the 25th has been cancelled due to industrial action on the part of AUT membership. MAR noted that the ROE site web pages have been reorganised and that many links in existing pages are now broken. This has not affected the WSA private area network (nor the old SSS/SHS server for that matter).